OpenStudy (anonymous):

All of the streets are depicted as either perpendicular or parallel lines. The equation of the lane passing through A and B is -7x + 3y = -21.5. What is the equation of the central street PQ?

OpenStudy (anonymous):

any help?

OpenStudy (anonymous):

-3x + 4y = 3 -1.5x − 3.5y = -31.5 2x + y = 20 -2.25x + y = -9.75

OpenStudy (texaschic101):

is PQ parallel or perpendicular to AB

OpenStudy (anonymous):

perpindicular

OpenStudy (texaschic101):

ok....perpendicular lines have negative reciprocal slopes. So first we find the slope of -7x + 3y = -21.5 3y = 7x - 21.5 -- divide by 3 y = 7/3x - 21.5/3 so the slope here is 7/3, but we need a negative reciprocal slope. To get that, we just flip the slope and change the sign. So the slope for the perpendicular line is -3/7 (see how I flipped the slope and changed the sign). -1.5x - 3.5y = -31.5 <--- this is your perpendicular equation -3.5y = 1.5x - 31.5 y = 1.5/-3.5x - 31.5/-3.5 y = - 3/7x + 9 <--- it has a negative reciprocal slope